OpenStudy (anonymous):

35 is what percent of 70?

OpenStudy (saifoo.khan):

Percent = % of = multiply so, \[\frac{35}{100} \times 70\]

OpenStudy (anonymous):

75%?

OpenStudy (saifoo.khan):

\[\frac{75}{100} \times anything\]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

35 is what percent of 70? 70x% = 35 -> x = 35/ .7 = 50%